bill was drawn on 14th June, 2006 at 8 months after date and was discounted on 24th September, 2006 at 5% per annum. If the banker’s gain on the basis of simple interest is Rs 3, calculate the sum for which the bill was drawn.
1
Expert's answer
2020-06-15T18:20:40-0400

A = Rs

i =5/100 =0.05

Bills drawn on 14th june so legally due date 14th feb

As bill is discounted on 26th sep

Remaining period =6 days of sep+31 days of oct + 30 days of nov +31 days of dec +31 days of jan + 16 days of feb


=145days


145/365 = 29/73 year


[3][0.05][49/73] =0.0596



The holder will receive = RS [3-0.0596]


= 2.9404